OpenStudy (supersquidsisters):

how would you solve 8(2 pi/3)? sorry im not too good with fractions and the pi is throwing me off

OpenStudy (google):

I don't understand how you wrote that expression

OpenStudy (supersquidsisters):

@Google its 8 times 2 pi over 3

OpenStudy (google):

8(2*pi /3)

OpenStudy (google):

\[8x \frac{ 2\pi }{ 3 }\]

OpenStudy (google):

?

OpenStudy (supersquidsisters):

@Google no not 8x just 8, other than that thats right

OpenStudy (google):

but its 8* 2pi/3 ?

OpenStudy (supersquidsisters):

@Google oh yeah it it 8 times, i thought you were using x as a variable sorry

OpenStudy (google):

Ok so let me help you

OpenStudy (google):

Lets break this problem down. 2pi = 2*pi so 2 x 3.14 = ?

OpenStudy (supersquidsisters):

@Google 6.28 ?

OpenStudy (google):

Yes be more confident with you math. ok so next part \[8*\ \frac{ 6.28 }{ 3 }\]

OpenStudy (google):

whats wrong with that problem?

OpenStudy (supersquidsisters):

@Google theres a decimal as the numerator

OpenStudy (google):

No.

OpenStudy (supersquidsisters):

(also sorry for the late reply my mom called )

OpenStudy (supersquidsisters):

@Google ok itf it isnt that then i have no idea im sorry

OpenStudy (google):

Its ok

OpenStudy (google):

\[8*(\frac{ 6.28 }{ 3 })\]

OpenStudy (supersquidsisters):

@Google oooooooh

OpenStudy (google):

Yeah

OpenStudy (google):

Is it multiple choice?

OpenStudy (supersquidsisters):

@Google yea my choices are 26 pi over 3 cm 10 pi over 3 cm 16 pi over 3 cm pi over 3 cm

OpenStudy (supersquidsisters):

i think its 16 pi over 3?

OpenStudy (google):

I am pretty sure it is

OpenStudy (supersquidsisters):

@Google oh ok, thank you!!

OpenStudy (google):

No problem